OpenStudy (anonymous):

Which function's graph has a period of 2? y=cos(x-pi/2) y=3cos pi x y=2 sin 3x y=-4 sin 2x

terenzreignz (terenzreignz):

The function takes the form \[\huge y=a\sin(bx+c)+d\]in general. By the way, it doesn't matter here, if sin is replaced by cos. The period is given by this formula \[\huge \frac{2\pi}{b}\]

OpenStudy (anonymous):

so it would be y= 2 sin 3x

jimthompson5910 (jim_thompson5910):

no it's not y = 2*sin(3x)

jimthompson5910 (jim_thompson5910):

that has a period of T = 2pi/B T = 2pi/3

OpenStudy (anonymous):

y=3cos pi x? 2 pi/pi

terenzreignz (terenzreignz):

Yes, that's right ^.^

terenzreignz (terenzreignz):

\[\Large \frac{2\pi}{\pi}=2\]

OpenStudy (anonymous):

thank you for explaining how to do it

OpenStudy (anonymous):

y= -4 sin pi x